Cullumber Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blankets retail for $45 throughout the country to loyal alumni of over 4,000 schools. Cullumber’s variable costs are 42% of sales; fixed costs are $116,000 per month.

Assume that variable costs increase to 45% of the current sales price and fixed costs increase by $12,500 per month. If Cullumber were to raise its sales price by 11% to cover these new costs, what would be the new annual breakeven point in sales dollars? (Round sales price to 2 decimal places, e.g. 52.75 and final answer to 0 decimal places, e.g. 5,275.)
Answer #1

Current position:

Selling price per unit = $45

Variable cost per unit = 42% of sales

= 45 x 42%

= $18.9

Fixed cost = $116,000

New Position:

Variable cost per unit = 45% of current selling price

= 45% x 45

= $20.25

Increase in fixed cost = $12,500

Hence, fixed cost = 116,000+12,500

= $128,500

Increase in selling price = 11%

= 45 x 11%

= $4.95

Hence, selling price = 45+4.95

= $49.95

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 49.95-20.25

= $29.70

Contribution margin ratio = Contribution margin per unit / Selling price per unit

= 29.70/49.95

= 59.4594595%

Break even point ( in dollar) = Fixed cost / Contribution margin ratio

= 128,500/59.4594595%

= $216,114

Kindly comment if you need further assistance. Thanks‼!